Circular No. 7893

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J0929–314 R. A. Remillard, Massachusetts Institute of Technology; and J. Swank

and T. Strohmayer, Goddard Space Flight Center, NASA, report the discovery of 185-Hz pulsations in XTE J0929–314 (IAUC 7889). This source was observed in a brief pointed observation with RXTE on May 2, while the average flux was 28 mCrab (2–30 keV). A power spectrum was computed for 800 s of PCA data, and a highly significant pulsation is seen at 185.09 Hz, with a strong harmonic at 370.18 Hz. This is the third known pulsar in which pulsations faster than 10 ms can be seen in the persistent x-ray emission. Rasters across the source give an improved position α = 9^{h}29^{m}18^{s}, δ = −31^{o}23′.1 (equinox J2000.0; systematic uncertainty 1′), consistent with the optical candidate (IAUC 7889).

M. P. Rupen, V. Dhawan, and A. J. Mioduszewski, National Radio Astronomy Observatory, report the detection of a radio counterpart to the x-ray transient XTE J0929–314 (IAUC 7888). Observations with the Very Large Array (VLA) at 4.86 GHz show a source with flux density 0.31 ± 0.07 mJy on May 3, and 0.36 ± 0.05 mJy on May 7, at α = 9^{h}29^{m}20^{s}.194, δ = −31^{o}23′03′′.41 (equinox J2000.0; uncertainty ± 0′′.3). This is 0′′.8 from the optical position reported by Greenhill et al. (IAUC 7889) and provides strong evidence that this optical identification is correct. Further optical and x-ray observations are strongly encouraged.

P. Cacella, Brasilia, Brazil, reports that an unfiltered CCD image taken with a 0.25-m reflector shows a variable (mag 18.3) that is possibly the optical counterpart to XTE J0929–314 at position end figures 20^{s}.22, 03′′.6.

SUPERNOVA 2002ck IN UGC 10030 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2002ck (cf. IAUC 7884), obtained by M. Calkins on May 2.45 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 8953 km/s for the host galaxy, the supernova expansion velocity is ∼ 10400 km/s for Si II (rest 635.5 nm) and 14400 km/s for Ca II (rest 395.1 nm). The spectral-feature age of the supernova (cf. IAUC 7774) is 1 ± 2 days before maximum light.

2002 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green